OpenStudy (anonymous):

A PAIR OF DICE IS ROLLED WHAT IS THE PROBABILITY OF ROLLING A SUM LESS THAN 6

OpenStudy (anonymous):

83%

OpenStudy (anonymous):

sum < 6 (1,1) (1,2) (1, 3) (1, 4) (2,3) (3,2) (4,1) (3,1) (2, 1) no. of possible events = 9 total number of events = 36 P(sum<6) = 9/36 = 1/4